OpenStudy (anonymous):

solve (5x-4)^1/2 = 6

OpenStudy (anonymous):

(5x-4)^(1/2) just means that its square root of (5x-4). you can square both sides, that will give you 5x-4= 36 5x= 36+4 x = (40)/(5) x = 8 hope thats the correct answer!

OpenStudy (anonymous):

Thanks!